# Bayesian Survival Analysis by Approximate Inference of Neural Networks

### 09 Apr 2024 v1

@@ Abstract
Predicting future events always comes with uncertainty, but traditional non-Bayesian methods cannot distinguish certain from uncertain predictions or explain the confidence in their predictions. In survival analysis, Bayesian methods applied to state-of-the-art solutions in the healthcare and biomedical field are still novel, and their implications have not been fully evaluated. In this paper, we study the benefits of modeling uncertainty in deep neural networks for survival analysis with a focus on prediction and calibration performance. For this, we present a Bayesian deep learning framework that consists of three Bayesian network architectures, which we train by optimizing the Cox partial likelihood and combining input-dependent aleatoric uncertainty with model-specific epistemic uncertainty. This enables us to provide uncertainty estimates as credible intervals when predicting the survival curve or as a probability density function over the predicted median survival times. For our empirical analyses, we evaluated our proposed method on four benchmark datasets and found that our method demonstrates prediction performance comparable to the state-of-the-art based on the concordance index and outperforms all other Cox-based approaches in terms of the mean absolute error. Our work explicitly compares the extent to which different Bayesian approximation techniques differ from each other and improves the prediction over traditional non-Bayesian alternatives.
@@ 1 Introduction
In this work, we propose three Bayesian architectures for survival analysis: Spectral-normalized Neural Gaussian Process (SNGP), Variational Inference (VI), and Monte Carlo Dropout (MCD). We provide a thorough evaluation of their performance in terms of prediction and calibration performance, and compare them to established literature benchmarks. The use of VI to train BNNs for survival analysis is popular in the literature [11, 15, 14], but requires doubling the number of network parameters. This motivates the use of variational approximation techniques to Bayesian uncertainty estimation, but also mandates an empirical evaluation to check if the approximate predictive distribution keeps the qualitative features of the exact distribution. A preliminary version of this work has been reported in [13]. The source code is available on GitHub [^1] https://github.com/thecml/baysurv.
@@ 3.4 Our proposed method
For the purpose of comparability in the results, consistency in the hyperparameters, and evaluation fairness, we propose a multilayer perceptron (MLP) backbone network architecture. An MLP is a fully connected feedforward NN; the number of hidden layers $L$ and neurons are hyperparameters tuned based on the data, see, e.g., [4]. The adoption of an MLP is aligned with the literature, as capable of providing solid predictive performance[36, 9, 42]. The proposed architecture is depicted in Fig. 1. There are three Bayesian configurations:
@@
    Bayesian approach trained with SNGP. The predictive (epistemic) uncertainty of SNGP is computed using the Laplace approximation and allows estimating the predictive uncertainty on a per-sample basis without MC sampling. Given a batch input $\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{N}\}$, the GP layer returns predictions $\{y_{1},y_{2},...,y_{N}\}$ and a covariance matrix $\Sigma^{N\times N}$, i.e., the predictive covariance matrix of the batch logits. The model is optimized according to the likelihood in Eq. (3.2). See Fig. 0(c).
@@
    Bayesian approach trained with VI, learning the variational approximation by optimizing Eq. (3.3), where weights and biases are treated as random variables. The output nodes provide the mean and standard deviation of the output $\hat{y}_{i}$ as a Gaussian draw, i.e., $\hat{y}_{i}\sim\mathcal{N}\left(\mu_{\bm{x}_{i}},\sigma_{\bm{x}_{i}}\right)$ for a $d$-dimensional sample $\bm{x}_{i}$. The Bayesian framework captures the epistemic uncertainty in the model, and sampling the output from a Gaussian captures the heteroscedastic aleatoric uncertainty. See Fig. 0(a).
@@
    Bayesian approach trained with MCD. We set dropout rates to $p=\{0.1,0.2,0.5\}$, capturing both epistemic and heteroscedastic aleatoric uncertainty. This corresponds to a deterministic MLP, but with Dropout at test time and optimized according to the likelihood function in Eq. 3.2. The output consists of two nodes corresponding to the mean and variance of the predictive Gaussian distribution, from which the $y_{i}$s are drawn conditionally on $\bm{x}_{i}$ and the random network configuration at each epoch due to the random deactivation of some of the nodes. See Fig. 0(b).
@@ 4.1 Setup
We adopt several evaluation metrics to assess the predictive and calibration performance of our baseline MLP model, its Bayesian variants, and literature benchmarks. To evaluate predictive performance, we report Antolini’s concordance index (CI) [52], the mean absolute error (MAE) using a hinge loss [53], the MAE using pseudo-observations [53] and the Integrated Brier Score (IBS) [54]. To evaluate calibration performance, we report the Integrated Calibration Index (ICI) [55], Distribution Calibration (D-Cal) [56] and Coverage Calibration (C-Cal) [14]. See Appendix 6 for more details on performance metrics.
@@ 4.2 Prediction results
Table 1 shows the predictive performance of our baseline MLP and its Bayesian variants, and literature benchmarks on the test sets. For VI and MCD, the reported metrics are constructed based on predictive means over 100 posterior draws. The proposed neural network architectures were trained for 100 epochs using early stopping with five epochs of patience. Concerning the baseline MLP model and the SNGP, VI and MCD variants, on the smallest METABRIC dataset (see Tab. 1(a)), all Bayesian models report a higher concordance index (CItd) and the MCD especially shows a significantly lower MAE (MAEH and MAEPO) than their baseline counterpart, thus modeling both aleatoric and epistemic uncertainty using VI or MCD improves predictive performance. In the mid-sized SEER dataset (see Tab. 1(b)), all Bayesian variants give approximately the same concordance index, but have a significantly lower MAEH than the baseline. On the contrary, in the larger SUPPORT dataset (see Tab. 1(c)), the SNGP and Bayesian variants struggle to keep up with the baseline MLP in all comparable metrics, with only MCD ($p$ = 0.5) outperforming both the baseline and its rival Bayesian models. In the large MIMIC-IV dataset (see Tab. 1(d)), the effect of the different estimation approaches is reduced, and all measures are generally aligned with each other, although VI and MCD still have a slight edge when it comes to MAEH. We observe that for all datasets, setting a dropout rate of $p=0.5$ when using MCD to approximate the VI solution leads to better convergence and better predictive performance than the baseline MLP.
@@ 4.3 Calibration results
Table 2: Calibration performance on the METABRIC, SEER, SUPPORT and MIMIC-IV test sets. Calibration is reported by a model’s $p$-value according to a Pearson’s $\chi^{2}$ test. Calibrated models have $p$-values greater than 0.05, indicated by Yes/No. A $p$-value close to one suggests good calibration. C-calibration is only applicable for models that can predict credible regions (CrIs) [14]. $N$: total sample size, $C$: pct. of censored data, $d$: number of covariates.
@@
Table 2 shows the calibration performance of our baseline MLP and its SNGP, VI, and MCD, as well as variants and literature benchmarks on the test sets. Regarding the baseline MLP model and its Bayesian variants, all models are D-calibrated on the small METABRIC dataset (see Tab. 2(a)), with the VI and MCD variants also being C-calibrated. We assess the agreement between predicted and observed survival probabilities using the ICI metric, i.e., the absolute difference between the calibration curve and the diagonal line of perfect calibration. We see that the MCD models improve on the baseline in terms of ICI and thus provide a better-calibrated survival curve. Using VI, despite being distribution and coverage calibrated, has a four-time higher ICI score than the MCD ($p$ = 0.5) in the METABRIC dataset. In the mid-sized SEER dataset (see Tab. 2(b)), all proposed architectures are again properly distributed and coverage calibrated; however, we do observe a slight decrease in calibration performance in terms of ICI when adopting the VI and MCD methods, which is also evident in the lower $p$-values for D-Cal. In the SUPPORT dataset (see Tab. 2(c)), only the SNGP variant is statistically D-calibrated when compared to the other Bayesian approaches, but their ICI scores are not markedly higher than the baseline or the SNGP. We observe a similar trend in the large MIMIC-IV dataset (see Tab. 2(d)), as none of the Bayesian methods are D-calibrated except the SNGP. Regarding the tight coupling between the quality of the MCD uncertainty estimates and the parameter choices, e.g., dropout layers, how many to use, and the choice of dropout rate, we conduct empirical experiments with various dropout rates to measure its effect. In terms of calibration performance, we do not see significant differences using either a low or high dropout rate; however, in the SEER, SUPPORT, and MIMIC-IV datasets, setting $p$ = 0.1 does provide higher $p$-values for D-calibration.
@@
Figure 3 shows the calibration curves of the baseline and its Bayesian variants in the METABRIC dataset visually. We compare the survival models against the observed frequencies of event at four percentiles (25th, 50th, 75th, and 90th) of the survival time and create a smoothed calibration curve using a flexible spline regression model. The ranges of the predicted probabilities in the test set at the 90th percentile are as follows: 0.239 to 1 (MLP), 0.413 to 0.998 (SNGP), 0.208 to 0.973 (VI), 0.261 to 1 (MCD, $p$ = 0.1), 0.086 to 1 (MCD, $p$ = 0.2) and 0.097 to 1 (MCD, $p$ = 0.5). While the calibration curves for the MCD and SNGP variants are visually better than the VI, it is still difficult to determine from the plot alone which model is better calibrated at various survival times.
Figure 2: Illustration of a workflow involving a model that is mindful of uncertainties. If a model prediction satisfies a certain confidence threshold, we trust it; otherwise, it is passed on to a human for additional evaluation.
Figure 3: Smoothed calibration curves for the baseline MLP model and its Bayesian variants in the METABRIC dataset. Left: observed versus predicted probability of event at four percentiles (25th, 50th, 75th and 90th) of the survival time. Right: kernel density plot visualizing the distribution of the predicted survival probabilities in the dataset.
@@
Figure 4: Model prediction in the METABRIC dataset using VI and MCD on a randomly selected sample from the test set. The prediction is made by drawing 1000 samples from the predictive distribution. Left column: prediction of the individual survival distribution (ISD) and the associated 90% credible interval (shaded region) with the observed event time. The solid black curve is the mean ISD, and the surrounding CrIs represent the uncertainty in the prediction. The predicted event time is the median of the survival curve (the cross point of the black curve and the vertical blue dashed line). Right column: probability density function of the predicted median survival times. Although the VI model makes a better prediction of the survival time ($\hat{y}_{i}=128$) than MCD ($\hat{y}_{i}=122$), the variance in its prediction is significantly higher. This shows the importance of communicating the underlying variance when making a prediction, even when the model provides a satisfactory answer on average.
@@ 4.4 Uncertainty estimation
We adopt the notion of credible intervals (CrIs) from [14]. CrIs in the Bayesian framework offer a way to quantify the uncertainty associated with a prediction. A credible interval is defined by two values, referred to as the lower and upper bounds, which enclose a certain percentage of the predictive distribution. The main difference between Bayesian CrIs and their frequentist counterpart, confidence intervals (CIs), is that CrIs treat the parameter of interest as a random variable and the bounds as fixed values, while CIs treat the parameter of interest as a fixed value and the bounds as random variables. Figure 4 shows two ways in which uncertainty can be expressed along with a survival prediction for decision support purposes. For this, we selected a random sample from the METABRIC test set and drew 1000 samples from the model’s predictive distribution to estimate the survival curve. Although both MCD (upper panel) and VI (lower panel) agree on the probability of event at the median (0.5) survival time, $\hat{y_{t}}$, the latter predicts a higher amount of variance around the survival curve and in the probability density function than the former. The width of the CrIs expresses the model’s uncertainty, which is especially useful when making predictions about novel individuals (new data samples). Adopting an AI model that can express uncertainty about its predictions can help medical practitioners make tailor-made treatment plans with an extra level of risk assessment and understanding to it. This in turn can increase the level of trust and confidence that practitioners, patients and relatives put in the prediction. Non-probabilistic models cannot offer these advantages, which limit their applicability in a setting, where the confidence in an outcome is important, e.g., clinical or medical practices.
@@ 5 Conclusion
Assessing uncertainty in high-risk domains, such as healthcare, is of utmost importance for boosting the point prediction that traditional approaches deliver. Figure 2 illustrates the workflow of using a model that can provide uncertainties along with its prediction. Given a confidence threshold, we can choose to trust its prediction or defer it to a human. This is the main advantage of using Bayesian methods, since traditional non-Bayesian methods cannot distinguish certain from uncertain predictions. In particular, in survival analysis, Bayesian methods applied to state-of-the-art solutions are very novel, and their implications are not extensively assessed. In this work, we have applied and evaluated three Bayesian architectures that allow to model and include aleatoric, epistemic, and both uncertainties for survival analysis. Over a wide set of traditional benchmark models, widely employed reference datasets, and existing models in the literature, we show the benefit of employing SNPG networks, Variational Inference (VI), and Monte Carlo Dropout (MCD) as effective solutions for modeling uncertainties in deep neural networks for survival analysis. Supported by a number of tables, detailed analyses, and graphical illustrations, our main takeaways are as follows:
  • The adoption of Bayesian techniques exhibits state-of-the-art predictive performance in real-world clinical datasets, and in particular, using MCD with a high dropout rate ($p$ = 0.5) provides better survival estimates in terms of MAE than all benchmark models, except MTLR. Our empirical experiments have shown that adopting the Bayesian framework does not worsen predictive performance compared to deterministic models, but actually improves them in many scenarios.
  • We find that in small-sized datasets, the goodness-of-fit test for D-calibration is easier satisfied for the Bayesian models compared to large-sized datasets. The baseline MLP and the Bayesian SNGP variant satisfy the assumption in all datasets, but introducing more degrees of freedom in the model and modeling heteroscedastic variance in the output, makes it easier to fail the hypothesis test. For example, the linear CoxPH model is almost perfectly calibrated in all scenarios, but has worse predictive performance than the Bayesian models.
@@
Our work is important because, for the first time, it explicitly compares the extent to which different Bayesian approximation techniques using deep learning differ from each other and improves the prediction over traditional non-Bayesian alternatives. Future research may extend this work towards the direction of the latest machine learning research in probabilistic learning, such as conformal prediction, or in survival prediction, such as recent frameworks that do not assume proportional hazards.

### 12 Apr 2024 v2

# Probabilistic Survival Analysis by Approximate Bayesian Inference of Neural Networks

@@ Abstract
Predicting future events always comes with uncertainty, but traditional non-probabilistic methods cannot distinguish certain from uncertain predictions. In survival analysis, probabilistic methods applied to state-of-the-art solutions in the healthcare and biomedical field are still novel, and their implications have not been fully evaluated. In this paper, we study the benefits of modeling uncertainty in deep neural networks for survival analysis with a focus on prediction and calibration performance. For this, we present a Bayesian deep learning framework that consists of three probabilistic network architectures, which we train by optimizing the Cox partial likelihood and combining input-dependent aleatoric uncertainty together with epistemic uncertainty. This enables us to provide uncertainty estimates as credible intervals when predicting the survival curve or as a probability density function over the predicted median survival times. For our empirical analyses, we evaluated our proposed method on four benchmark datasets and found that our method demonstrates prediction performance comparable to the state-of-the-art based on the concordance index and outperforms all other Cox-based approaches in terms of the mean absolute error. Our work explicitly compares the extent to which different Bayesian approximation techniques differ from each other and improves the prediction over traditional non-probabilistic alternatives.
@@ 1 Introduction
In this work, we propose three probabilistic architectures for survival analysis: Spectral-normalized Neural Gaussian Process (SNGP), Variational Inference (VI), and Monte Carlo Dropout (MCD). We provide a thorough evaluation of their performance in terms of prediction and calibration, and compare them to established literature benchmarks. The use of VI to train BNNs for survival analysis is popular in the literature [11, 15, 14], but requires doubling the number of network parameters. This motivates the use of variational approximation techniques to Bayesian uncertainty estimation, but also mandates an empirical evaluation to check if the approximate predictive distribution keeps the qualitative features of the exact distribution. A preliminary version of this work has been reported in [13]. The source code is available on GitHub [^1] https://github.com/thecml/baysurv.
@@ 3.4 Our proposed method
For the purpose of comparability in the results, consistency in the hyperparameters, and evaluation fairness, we propose a multilayer perceptron (MLP) backbone network architecture. An MLP is a fully connected feedforward NN; the number of hidden layers $L$ and neurons are hyperparameters tuned based on the data, see, e.g., [4]. The adoption of an MLP is aligned with the literature, as capable of providing solid predictive performance[36, 9, 42]. The proposed architecture is depicted in Fig. 1. There are three probabilistic configurations:
@@
    Probabilistic approach trained with SNGP. The predictive (epistemic) uncertainty of SNGP is computed using the Laplace approximation and allows estimating the predictive uncertainty on a per-sample basis without MC sampling. Given a batch input $\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{N}\}$, the GP layer returns predictions $\{y_{1},y_{2},...,y_{N}\}$ and a covariance matrix $\Sigma^{N\times N}$, i.e., the predictive covariance matrix of the batch logits. The model is optimized according to the likelihood in Eq. (3.2). See Fig. 0(c).
@@
    Probabilistic approach trained with VI, learning the variational approximation by optimizing Eq. (3.3), where weights and biases are treated as random variables. The output nodes provide the mean and standard deviation of the output $\hat{y}_{i}$ as a Gaussian draw, i.e., $\hat{y}_{i}\sim\mathcal{N}\left(\mu_{\bm{x}_{i}},\sigma_{\bm{x}_{i}}\right)$ for a $d$-dimensional sample $\bm{x}_{i}$. The Bayesian framework captures the epistemic uncertainty in the model, and sampling the output from a Gaussian captures the heteroscedastic aleatoric uncertainty. See Fig. 0(a).
@@
    Probabilistic approach trained with MCD. We set dropout rates to $p=\{0.1,0.2,0.5\}$, capturing both epistemic and heteroscedastic aleatoric uncertainty. This corresponds to a deterministic MLP, but with MC dropout at testinference time and optimized according to the likelihood function in Eq. 3.2, and two output nodes corresponding to the mean and variance of the predictive Gaussian distribution, from which the $y_{i}$s are drawn conditionally on $\bm{x}_{i}$ and the random network configuration at each epoch due to the random deactivation of some of the nodes. See Fig. 0(b).
@@ 4.1 Setup
We adopt several evaluation metrics to assess the predictive and calibration performance of our baseline MLP model, its probabilistic variants, and literature benchmarks. To evaluate predictive performance, we report Antolini’s concordance index (CI) [52], the mean absolute error (MAE) using a hinge loss [53], the MAE using pseudo-observations [53] and the Integrated Brier Score (IBS) [54]. To evaluate calibration performance, we report the Integrated Calibration Index (ICI) [55], Distribution Calibration (D-Cal) [56] and Coverage Calibration (C-Cal) [14]. See Appendix 6 for more details on performance metrics.
@@ 4.2 Prediction results
Table 1 shows the predictive performance of our baseline MLP and its probabilistic variants, and literature benchmarks on the test sets. For VI and MCD, the reported metrics are constructed based on predictive means over 100 posterior draws. The proposed neural network architectures were trained for 100 epochs using early stopping with five epochs of patience. Concerning the baseline MLP model and the SNGP, VI and MCD variants, on the smallest METABRIC dataset (see Tab. 1(a)), all probabilistic models report a higher concordance index (CItd) and the MCD especially shows a significantly lower MAE (MAEH and MAEPO) than their baseline counterpart, thus modeling both aleatoric and epistemic uncertainty using VI or MCD improves predictive performance. In the mid-sized SEER dataset (see Tab. 1(b)), all probabilistic variants give approximately the same concordance index, but have a significantly lower MAEH than the baseline. On the contrary, in the larger SUPPORT dataset (see Tab. 1(c)), the SNGP and Bayesian variants struggle to keep up with the baseline MLP in all comparable metrics, with only MCD ($p$ = 0.5) outperforming both the baseline and its rival probabilistic models. In the large MIMIC-IV dataset (see Tab. 1(d)), the effect of the different estimation approaches is reduced, and all measures are generally aligned with each other, although VI and MCD still have a slight edge when it comes to MAEH. We observe that for all datasets, setting a dropout rate of $p=0.5$ when using MCD to approximate the VI solution leads to better convergence and better predictive performance than the baseline MLP.
@@ 4.3 Calibration results
Table 2: Calibration performance on the METABRIC, SEER, SUPPORT and MIMIC-IV test sets. D/C-Calibration and C-calibration areis reported by a model’s $p$-value according to a Pearson’s $\chi^{2}$ goodness-of-fit test. Calibrated models have $p$-values greater than 0.05, indicated by a Yes/No. A $p$-value close to one suggests good calibration. A $p$-value close to one suggests good calibration. C-calibration is only applicable for models that can predict credible regions (CrIs) [14]. $N$: total sample size, $C$: pct. of censored data, $d$: number of covariates.
@@
Table 2 shows the calibration performance of our baseline MLP and its SNGP, VI, and MCD, as well as variants and literature benchmarks on the test sets. Regarding the baseline MLP model and its probabilistic variants, all models are D-calibrated on the small METABRIC dataset (see Tab. 2(a)), with the VI and MCD variants also being C-calibrated. We assess the agreement between predicted and observed survival probabilities using the ICI metric, i.e., the absolute difference between the calibration curve and the diagonal line of perfect calibration. We see that the MCD models improve on the baseline in terms of ICI and thus provide a better-calibrated survival curve. Using VI, despite being distribution and coverage calibrated, has a four-time higher ICI score than the MCD ($p$ = 0.5) in the METABRIC dataset. In the mid-sized SEER dataset (see Tab. 2(b)), all proposed architectures are again properly distributed and coverage calibrated; however, we do observe a slight decrease in calibration performance in terms of ICI when adopting the VI and MCD methods, which is also evident in the lower $p$-values for D-Cal. In the SUPPORT dataset (see Tab. 2(c)), only the SNGP variant is statistically D-calibrated when compared to the other probabilistic approaches, but their ICI scores are not markedly higher than the baseline or the SNGP. We observe a similar trend in the large MIMIC-IV dataset (see Tab. 2(d)), as none of the probabilistic methods are D-calibrated except the SNGP. Regarding the tight coupling between the quality of the MCD uncertainty estimates and the parameter choices, e.g., dropout layers, how many to use, and the choice of dropout rate, we conduct empirical experiments with various dropout rates to measure its effect. In terms of calibration performance, we do not see significant differences using either a low or high dropout rate; however, in the SEER, SUPPORT, and MIMIC-IV datasets, setting $p$ = 0.1 does provide higher $p$-values for D-calibration.
@@
Figure 3 shows the calibration curves of the baseline and its probabilistic variants in the METABRIC dataset visually. We compare the survival models against the observed frequencies of event at four percentiles (25th, 50th, 75th, and 90th) of the survival time and create a smoothed calibration curve using a flexible spline regression model. The ranges of the predicted probabilities in the test set at the 90th percentile are as follows: 0.239 to 1 (MLP), 0.413 to 0.998 (SNGP), 0.208 to 0.973 (VI), 0.261 to 1 (MCD, $p$ = 0.1), 0.086 to 1 (MCD, $p$ = 0.2) and 0.097 to 1 (MCD, $p$ = 0.5). While the calibration curves for the MCD and SNGP variants are visually better than the VI, it is still difficult to determine from the plot alone which model is better calibrated at various survival times.
Figure 2: Illustration of a workflow involving a model that is mindful of uncertainties. If the model prediction satisfies a confidence threshold, we trust it; otherwise, it is passed on to a human for additional evaluation.
Figure 3: Smoothed calibration curves for the baseline MLP model and its probabilistic variants in the METABRIC dataset. Left: observed versus predicted probability of event at four percentiles (25th, 50th, 75th and 90th) of the survival time. Right: kernel density plot visualizing the distribution of the predicted survival probabilities in the dataset.
@@
Figure 4: Model prediction in the METABRIC dataset using the VI and MCD model on a randomly selected sample from the test set. The prediction is made by drawing 1000 samples from the predictive distribution. Left column: prediction of the individual survival distribution (ISD) and the associated 90% credible interval (shaded region) with the observed event time. The solid black curve is the mean ISD, and the surrounding CrIs represent the uncertainty in the prediction. The predicted event time is the median of the survival curve (the cross point of the black curve and the vertical blue dashed line). Right column: probability density function of the predicted median survival times. Although the VI model makes a better prediction of the survival time ($\hat{y}_{i}=128$) than MCD ($\hat{y}_{i}=122$), the variance in its prediction is significantly higher. This shows the importance of communicating the underlying variance when making a prediction, even when the model provides a satisfactory answer on average.
@@ 4.4 Uncertainty estimation
We adopt the notion of credible intervals (CrIs) from [14]. CrIs in the Bayesian framework offer a way to quantify the uncertainty associated with a prediction. A credible interval is defined by two values, referred to as the lower and upper bounds, which enclose a certain percentage of the predictive distribution. The main difference between Bayesian CrIs and their frequentist counterpart, confidence intervals (CIs), is that CrIs treat the parameter of interest as a random variable and the bounds as fixed values, while CIs treat the parameter of interest as a fixed value and the bounds as random variables. Figure 4 shows two ways in which uncertainty can be expressed along with a survival prediction for decision support purposes. For this, we selected a random sample from the METABRIC test set and drew 1000 samples from the predictive distribution of the model to estimate the survival curve. Although both MCD (upper panel) and VI (lower panel) agree on the median survival time, $\hat{y_{t}}$, the latter predicts a higher amount of variance around the survival curve and in the probability density function than the former. The width of the CrIs expresses the model’s uncertainty, which is especially useful when making predictions about novel individuals (new data samples). Adopting an AI model that can express uncertainty about its predictions can help medical practitioners make tailor-made treatment plans with an extra level of risk assessment and understanding to it. This in turn can increase the level of trust and confidence that practitioners, patients and relatives put in the prediction. Non-probabilistic models cannot offer these advantages.
@@ 5 Conclusion
Assessing uncertainty in high-risk domains, such as healthcare, is of utmost importance for boosting the point prediction that traditional approaches deliver. Figure 2 illustrates the workflow of using a model that can provide uncertainties along with its prediction. Given a confidence threshold, we can choose to trust its prediction or defer it to a human. This is the main advantage of using probabilistic methods, since traditional non-probabilistic methods cannot distinguish certain from uncertain predictions. In particular, in survival analysis, probabilistic methods applied to state-of-the-art solutions are very novel, and their implications are not extensively assessed. In this work, we have applied and evaluated three probabilistic architectures that allow to model and include aleatoric, epistemic, and both uncertainties for survival analysis. Over a wide set of traditional benchmark models, widely employed reference datasets, and existing models in the literature, we show the benefit of employing SNPG networks, Variational Inference (VI), and Monte Carlo Dropout (MCD) as effective solutions for modeling uncertainties in deep neural networks for survival analysis. Supported by a number of tables, detailed analyses, and graphical illustrations, our main takeaways are as follows:
  • The adoption of probabilistic techniques exhibits state-of-the-art predictive performance in real-world clinical datasets, and in particular, using MCD with a high dropout rate ($p$ = 0.5) provides better survival estimates in terms of mean absolute error (MAE) than all benchmark models, except MTLR. Our empirical experiments have shown that adopting the Bayesian framework does not worsen predictive performance compared to deterministic models, but actually improves them in many scenarios.
  • We find that in small-sized datasets, the goodness-of-fit test for D-calibration is easier satisfied for the Bayesian models compared to large-sized datasets. The baseline MLP and the probabilistic SNGP variant satisfy the assumption in all datasets, but introducing more degrees of freedom in the model and modeling heteroscedastic variance in the output, makes it easier to fail the hypothesis test. For example, the linear CoxPH model is almost perfectly calibrated in all scenarios, but has worse predictive performance than the Bayesian models.
@@
Our work is important because, for the first time, it explicitly compares the extent to which different Bayesian approximation techniques using deep learning differ from each other and improves the prediction over traditional non-probabilistic alternatives. Future research may extend this work towards the direction of the latest machine learning research in probabilistic learning, such as conformal prediction, or recent frameworks that do not assume proportional hazards.